NERSCPowering Scientific Discovery Since 1974

January 19

PDSF Users Meeting 1/19/10

Attending: Eric and Iwona from PDSF and users Andrew, Andrei, Marjorie, Craig, Jeff A. and Jeff P.

Cluster status: Utilization has been light. STAR mentioned that they are limited by the io resources being set as low as they are and raising them was discussed.

Outages: Discussed the recent issues with pdsf1 and pdsf2.

Upcoming downtimes: None.

New hardware: Data transfer node is being shipped. eliza11 is going away so ATLAS is buying a replacement. This turned into a discussion about using SL5. Action items are getting it set up to run batch jobs (transparent to user based on chos setting) and documenting it on the web. pdsf1 is now SL5 along with pdsf5 and the other interactives will be upgraded in the future.

ATLAS grid node: Ready for ATLAS to test.

Squid/fuse: Squid is working well in Yushu's office for the time being.

Other Topics:

- STAR project accounts testing: pdsf5 has been set up and the testing is in progress.